Jack has 93 coins. Some are nickels, and some are quarters. He has 16.85. How many of each does he have?
Urgent, please assist.

Answer by checkley79(3341)

Q+N=93
Q=93-N
.25Q+.05N=16.85
.25(93-N)+.05N=16.85
23.25-.25N+.05N=16.85
-.20N=16.85-23.25
-.20N=-6.40
N=-6.40/-.20
N=32 NICKLES.
Q=93-32=61 QUARTERS.
PROOF:
.25*61+.05*32=16.85
15.25+1.60=16.85
16.85=16.85